# Foundation models for high-energy physics

### v1

@@ 2 Motivation
The remainder of this review will mainly concern works based on the third approach, with a focus on collider physics. Collider experiments at the LHC like ATLAS [15] and CMS [16] collect vast amounts of highly complex and structured data. The data, coming from different subdetectors, is inherently multimodal. In addition, simulated (and therefore, labeled) datasets of high quality exist. As foundation models need large amounts of data for pretraining, this setting provides an excellent opportunity to develop and implement such models, investigating their capabilities in capturing intricacies and correlations in the data that we would perhaps not be able to untangle otherwise. Furthermore, physicists perform a wide range of downstream tasks using this data, including event classification, object tagging, anomaly detection, clustering, generation and regression. Machine learning methods are already being implemented in several of the experimental stages, from triggering [17, 18] to analysis [19, 20], meaning that expertise and experience already exist in the field. Having access to a rich latent representation of the data could also boost the achievable performance on small datasets, relevant in the context of rare processes where the amount of available simulation is limited due to narrow selection criteria.
@@ 4 Comparing existing foundation models
This section includes models that fit the foundation model criteria as outlined in the introduction, and focus on collider physics including particle jets [^3] Foundation models for collider physics not including particle jets include applications to nuclear physics [44, 45].. Some of these models do not refer to themselves as foundation models, but since they fit the criteria they are included nonetheless to show what different approaches have been investigated so far [^4] Works that use similar techniques as the ones described in the following but do not perform any finetuning or indicate any foundation model intent, will not be included here..
  • Particle Transformer (ParT) [26] (Feb 2022) pretrains on a supervised classification task, utilizing jet constituent features: kinematic, particle ID, trajectory discplacement and interaction features (the latter inspired by the jet clustering history [46]). Downstream tasks include finetuning to classification on different datasets.
@@
  • Re-simulation-based self-supervised learning (RS3L) [48] (Mar 2024) pretrains using contrastive learning, aiming to group augmentations of the same simulated process and separate them from the other events. The augmentations are created by fixing the hard process and then re-running the showering, hadronization and detector response steps of the simulation [^5] See e.g. [49] for an introduction to the first three simulation steps at hadron colliders, and [50, 51] for a examples of detector response simulation methods.. Features used are jet constituent kinematic features, particle ID and trajectory displacement features. The model is finetuned to supervised classification.

### v2

@@ 2 Motivation
The remainder of this review will mainly concern works based on the third approach, with a focus on collider physics. Collider experiments at the LHC like ATLAS [15] and CMS [16] collect vast amounts of highly complex and structured data. The data, coming from different subdetectors, is inherently multimodal. In addition, simulated (and therefore, labeled) datasets of high quality exist. As foundation models need large amounts of data for pretraining, this setting provides an excellent opportunity to develop and implement such models, investigating their capabilities in capturing intricacies and correlations in the data that we would perhaps not be able to untangle otherwise. Furthermore, physicists perform a wide range of downstream tasks using this data, including event classification, object tagging, anomaly detection, clustering, generation and regression. Machine learning methods are already being implemented in several of the experimental stages, from triggering to analysis (see [17, 18, 19, 20] for a few examples), meaning that expertise and experience already exist in the field. Having access to a rich latent representation of the data could also boost the achievable performance on small datasets, relevant in the context of rare processes where the amount of available simulation is limited due to narrow selection criteria.
@@ 4 Comparing existing foundation models
This section includes models that fit the foundation model criteria as outlined in the introduction, and focus on collider physics including particle jets [^3] Foundation models for collider physics not including particle jets include nuclear physics [44, 45] applications.. Some of these models do not refer to themselves as foundation models, but since they fit the criteria they are included nonetheless to show what different approaches have been investigated so far [^4] Works that use similar techniques as the ones described in the following but do not perform any finetuning or indicate any foundation model intent, will not be included here..
  • Particle Transformer (ParT) [26] (Feb 2022) pretrains on a supervised classification task, utilizing jet constituent features: kinematic, particle ID, trajectory discplacement and interaction features (the latter inspired by the jet clustering history [46]). Downstream tasks include finetuning to classification on different datasets555[47] extended this to event classification by jointly optimizing the reconstruction (tagging) and analysis tasks..
@@
  • Re-simulation-based self-supervised learning (RS3L) [49] (Mar 2024) pretrains using contrastive learning, aiming to group augmentations of the same simulated process and separate them from the other events. The augmentations are created by fixing the hard process and then re-running the showering, hadronization and detector response steps of the simulation [^6] See e.g. [50] for an introduction to the first three simulation steps at hadron colliders, and [51, 52] for examples of detector response simulation methods.. Features used are jet constituent kinematic features, particle ID and trajectory displacement features. The model is finetuned to supervised classification.